How much heat energy is required to raise the temperature of 0.365 kg of copper from 23.0 ∘C to 60.0 ∘C? The specific heat of copper is 0.0920 cal/(g⋅∘C) .
1
Expert's answer
2018-01-16T14:19:06-0500
Q = cm∆T = cm(T2 – T1)
Q = 0.0920 * 365 * (60 -23) = 1242.46 cal

Answer:
1242.46 cal